OpenStudy (anonymous):

4. Write the equation of a line in point slope form that has a slope of 4 3 − and passes through the point (4, 5).

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

its suppose to be -3/4

OpenStudy (anonymous):

slope of what??? 4 or 3 or 4/3

OpenStudy (anonymous):

the slope of -3/4

OpenStudy (anonymous):

y-5=-3/4(x-4) y-5=-3x/4+3 y=-3x/4+8

OpenStudy (anonymous):

So the real question is: Write the equation of a line in point slope form that has a slope of -3/4 and passes through the point (4, 5).

OpenStudy (anonymous):